OpenStudy (anonymous):

(-5x^8power) (-7x^6power) is

OpenStudy (mathstudent55):

First, multiply the numbers: (-5)(-7) = 35 Then (x^8)(x^6) = x^14 Answer: 35x^14

OpenStudy (anonymous):

\[(-5x^8) (-7x^6) = (-5)(-7)x^{8+6}=...\]